### v1

@@ 4.6 Complexity Analysis
Algorithm 2 can be implemented in time $\tilde{O}(n^{6}W\log^{O(1)}(n))$ with the same level of guarantee for additive error and privacy.
Recall that we assume the bound of the sum of edge weights to be polynomial. Thus Algorithm 2 runs in polynomial time. The proof of the lemma is similar to EKKL (20), so here we only give a proof sketch.
@@
According to Theorem 40 and Lemma 41, we have $T=\Theta(\frac{W({\varepsilon}{U}_{\triangle}+{U}_{\Lambda})}{n\log(n/\delta)\tilde{\ell}_{3}(\hat{G})})=\tilde{O}(W)$. In each inner iteration of Algorithm 2, note that Algorithm 3 runs in $O(n^{2})$ time. Denote $\mathbf{X}^{*}$ be the maximizer of the SDP $\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X})$. We use the algorithm of LSW (15) to find an approximate solution of the SDP in Line 6 of Algorithm 2 in time $\tilde{O}(n^{6}\log^{O(1)}(n/\mu))$, such that, $\left\|(\mathbf{X}^{*})^{-\frac{1}{2}}(\mathbf{X}^{*}-\mathbf{X})(\mathbf{X}^{*})^{-\frac{1}{2}}\right\|_{F}\leq\mu$. Therefore, within $LT$ iterations, Algorithm 2 can be implemented in time $\tilde{O}(n^{6}W\log^{O(1)}(n))$. According to EKKL (20), the additive error only differs by a constant factor when we choose $\mu=1/n^{O(1)}$. And the privacy loss in each iteration of Algorithm 2 is still $O(\frac{\ell_{3}(\hat{G})}{\lambda})$, hence the privacy guarantee is the same as before. ∎
@@ 5 The Lower Bound
Given graph $G$, let $\ell_{h}(G)=\max_{(i,j)\in\binom{V}{2}}\sum_{V^{\prime}=(i,j,v_{1},v_{2},\dots,v_{h-2})\in\binom{V}{h}}\prod_{(k,\ell)\in\binom{V^{\prime}}{2}\setminus\{(i,j)\}}\mathbf{w}_{(k,\ell)}$ denotes the local sensitivity of $K_{h}$-motif cuts of $G$. Note that, its local sensitivity is defined as the maximum $K_{h}$-motif cut difference between $G$ and its neighboring graphs. For unweighted graph with maximum degree at most $d$, we have $\ell_{h}(G)=O(d^{h-2})$. For unweighted dense graph, we have $\ell_{h}(G)=\Theta(n^{h-2})$. For unweighted graph generated from $G(n,p)$, we have $\ell_{h}(G)=\Theta(n^{h-2}p^{\frac{h^{2}-h-2}{2}})$.
We show the following lower bound. Note that Theorem 4 follows from the following theorem by setting $h=3$ and $\beta=\frac{3}{4}$.
@@
Note that $\mathbf{A}$ is fixed and does not depend on $G$. Let $\mathbf{x}\in\{0,1\}^{\binom{n}{2}}$ be the indicator vector of $E$. Let $\mathbf{x}_{K_{h}}=f_{K_{h}}(x)\in\{0,1\}^{\binom{n}{h}}$ be the indicator vector of $K_{h}$ in $G$, i.e. for each tuple $i_{1},i_{2},\dots,i_{h}$ of $h$ different indices, $(\mathbf{x}_{K_{h}})_{i_{1},i_{2},\dots,i_{h}}=1$ if the subgraph induced by $i_{1},i_{2},\dots,i_{h}$ is $K_{h}$ and 00 otherwise. Then $\mathbf{A}\mathbf{x}_{K_{h}}$ specifies the $K_{h}$-motif size of all $(S,T)$-cuts in $G$, i.e., we have
$(\mathbf{A}\mathbf{x}_{K_{h}})_{(S,T)}=\sum_{I\in\mathcal{M}(G,K_{h})}\mathbf{1}_{\text{$I$ crosses $(S,T)$}}.$
@@
    for each vertex $v\in V$, its degree belongs to the interval $[\gamma n/2,2\gamma n]$;
@@
    for each edge $e\in E\cup E^{\prime}$, the number of $K_{h}$-instances containing $e$ belongs to the interval $[\frac{\ell_{h}(G)}{2},2\ell_{h}(G)]$;
@@
For any $i\leq h$ and any set $W\subseteq V$, we let $W^{(i)}$ denote the set of all subsets $R\subseteq W$ with exactly $i$ distinct vertices. For any $i\leq h$ and a set $R\subseteq V^{(i)}$, and $h-i$ other vertices $t_{i+1},\dots,t_{h}$, we let $\chi_{R,t_{i+1},\dots,t_{h}}$ denote the coloring of the potential $K_{h}$ instance induced by the vertex set $R\cup\{t_{i+1},\dots,t_{h}\}$. We define $\chi_{R}=\sum_{t_{i+1},\dots,t_{h}\in V^{(h-i)}}\chi_{R,t_{i+1},\dots,t_{h}}$ and define $\chi_{R,t}=\sum_{t_{i+2},\dots,t_{h}\in V^{(h-i-1)}}\chi_{R,t,t_{i+2},\dots,t_{h}}$. We extend the definition of $\chi$ to any tuple $i_{1},\dots,i_{h}\in V^{(h)}$ by setting $\chi_{i_{1},\dots,i_{h}}=0$ whenever there exists a repeated index in the tuple.
@@
Let $H=(W,E)$ be a graph such that for any fixed $B$ with $|B|=i-1$, the number of $t\in W$ such that $t$ is connected to all vertices in $B$ is at most $2n\gamma^{i-1}$. Let $S\cup T$ be a random partition of $W$, i.e., each vertex is independently assigned to each class with probability $1/2$. Then
@@
$\geq 2^{-4}\cdot\mathbb{E}\left[\sum_{R\in T_{2}^{(2)}}\left|y_{2}(R)\right|/\sqrt{n\gamma}$
@@
=$\mathcal{F}=\{\mathcal{J}\subseteq\{2,\dots,h\}:\left|\chi_{\mathcal{J}}\right|\geq(2h)^{-h+\left|\mathcal{J}\right|}\cdot\left|\chi_{S_{1}^{+},S_{2},\dots,S_{h}}\right|\}.$
@@
Let $\mathbf{x}$ be an indicator vector of the edge set of a graph $G=(V,E)$ which satisfies the following properties ()($\star$)( ⋆ ):
@@
    for each vertex $v\in V$, its degree belongs to the interval $[\gamma n/2,2\gamma n]$;
@@
    for each edge $e\in E$, the number of $K_{h}$-instances containing $e$ belongs to the interval $[\frac{\ell_{h}(G)}{2},2\ell_{h}(G)]$;
@@
Let $\mathcal{A}$ simply be the algorithm that output an indicator vector $x^{\prime}$ of any graph that satisfies the properties ()($\star$)( ⋆ ) and that
@@
Part 2) follows from the same argument as those in the proof of Lemma 5.3 in EKKL (20). For part 1), we describe our changes. Formally, we say that $\mathbf{x}\sim X_{\mid Y=\mathbf{y}}$ is good, if $\left\|\mathbf{A}\mathbf{x}_{K_{h}}-\mathbf{y}\right\|_{\infty}\leq\mathrm{disc}_{\mathcal{C}_{\sigma,\gamma}}(\mathbf{A})/2-1$ and the properties ()($\star$)( ⋆ ) given in the statement of Lemma 50 are satisfied.
Note that the property that $\mathbf{x}\sim X_{\mid Y=\mathbf{y}}$ is good is at least $(1-\frac{1}{\mathrm{poly}(n)})\cdot\beta$. This is true as our assumption that $G\sim G(n,p)$ and that $(\frac{\log n}{n})^{1/(h-1)}\ll p\leq\frac{1}{2}$, it holds that $G$ satisfies the properties ()($\star$)( ⋆ ) with probability at least $1-\frac{1}{\mathrm{poly}(n)}$. This then finishes the proof of the lemma. ∎
Finally, our lower bound Theorem 44 follows from the following lemma from EKKL (20), which in turn is built from Lemma 2.1.2 in Bun (16).
@@
Now consider the scaled version of $G$ described in Theorem 44, where the weight vector $\mathbf{w}$ of $G$ is scaled by a factor of $\frac{1}{\varepsilon}$, resulting in a new weight vector $\frac{1}{\varepsilon}\cdot\mathbf{w}$. Note that the total weight is $W=\frac{m}{\varepsilon}$. When we let $\overline{T}_{\mathbf{w}}=\sqrt{Wn}\cdot\ell_{h}(G)$, then $\overline{T}_{\frac{1}{\varepsilon}\cdot\mathbf{w}}=\frac{1}{\varepsilon^{\binom{h}{2}-\frac{1}{2}}}\overline{T}_{\mathbf{w}}$. By a similar analysis to Lemma 53, we can establish the bound
@@ A.1 Synthetic Graphs with Multiplicative Errors and Interactive Solutions
Finally, we note that an interactive solution for preserving privacy in the motif cut structure could also be considered (see Section 1.2 for considerations related to edge cuts). In this scenario, data analysts could specify any cut $(S,V\setminus S)$ with the aim of determining (with some acceptable error) the number of triangles or motifs connecting the two groups while maintaining privacy. However, our primary focus is on a stronger, non-interactive solution, i.e. to release a private synthetic dataset: a new, private graph that approximately preserves the motif cut function of the original graph.
@@ A.2 Some Tempting Approaches That Do Not Work
The third approach is to make use of the triangle-motif-weighted graph associated with the input graph [6]. That is, one can first convert the original graph ${G}$ into a triangle-motif weighted graph ${G}_{\triangle}$ with a weight vector $\mathbf{{w}}_{\triangle}$, where $\mathbf{{w}}_{\triangle}(e)$ denotes the sum of the weights of triangles containing the endpoints of edge $e$ simultaneously. For triangle-motif weighted graphs, there is a useful property: the size of the edge cut $(S,V\setminus S)$ in ${G}_{\triangle}$ is exactly twice the sum of the weights of triangles crossing the cut in $G$. Therefore, a naive approach would be to use a private edge cut release algorithm, such as the one proposed in [37], on the triangle-motif weighted graph. However, after applying the DP algorithm on ${G}_{\triangle}$, the resulting graph $H$ may not correspond to a triangle-motif weighted graph, i.e. there does not exist a graph whose triangle-motif weighted graph is $H$. As a result, through this approach, we can only privately release the values of triangle-motif cuts but cannot release a synthetic graph.
Appendix B Deferred Preprocessing Step in Section 3
@@
    Set $u_{\max}=\max_{e\in\binom{V}{2}}\mathbf{u}_{e}$. Recall that $\ell_{3}(\hat{G})=\max_{(i,j)\in\binom{V}{2}}\sum_{s\in V\setminus\{i,j\}}\mathbf{\hat{w}}_{(i,s)}\mathbf{\hat{w}}_{(j,s)}$ denotes the local sensitivity of triangle-motif cuts of $\hat{G}$. We further define ~$\tilde{\ell}_{3}(\hat{G})=\ell_{3}(\hat{G})+u_{\max}(\mathrm{Lap}(1/\varepsilon_{3})+\log(6n^{2}/\beta)/\varepsilon_{3})$.
By Lemma 10 and Lemma 6, the released $W$, $\mathbf{u}_{e}$ and ~$\tilde{\ell}_{3}(\hat{G})$ are $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$-differentially private respectively. By Lemma 11 and union bound, with probability at least $31-\frac{\beta}{3}-\binom{n}{2}\frac{\beta}{3n^{2}}\geq 1-2\beta/$ 3, it holds that
@@
Assume that $\hat{W}=\Omega(\frac{1/\beta}{\varepsilon})$ and $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$. We refer to this as the non-degenerate case. (When this assumption does not hold, we call it the degenerate case. We will provide a case analysis in Corollary 42.) Recall that $\ell_{3}(\overline{G})=\max_{(i,j)\in\binom{V}{2}}\sum_{s\in V\setminus\{i,j\}}\overline{\mathbf{w}}_{(i,s)}\overline{\mathbf{w}}_{(j,s)}$ is the local sensitivity of triangle cuts in $\overline{G}$. Since $\hat{W}=\Omega(\frac{1/\beta}{\varepsilon})$, we have $\ell_{3}(\overline{G})=(\frac{\overline{W}}{\hat{W}})^{2}\ell_{3}(\hat{G})=\Theta(\ell_{3}(\hat{G}))$. Additionally, by the assumption $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$ and the fact ~$\tilde{\ell}_{3}(\hat{G})\leq\ell_{3}(\hat{G})+2u_{\max}\log(6n^{2}/\beta)/\varepsilon_{3}\leq\ell_{3}(\hat{G})+O(w_{\max}\log^{2}(n/\beta)/\varepsilon_{2}\varepsilon_{3})$, $\varepsilon_{2}=\varepsilon_{3}=\varepsilon/6$, it holds that ~$\tilde{\ell}_{3}(\hat{G})=\Theta(\ell_{3}(\hat{G}))$.

### v2

@@ 4.6 Complexity Analysis
Algorithm˜2 can be implemented in time $\tilde{O}(n^{6}W\log^{O(1)}(n))$ with the same level of guarantee for additive error and privacy.
Recall that we assume the bound of the sum of edge weights to be polynomial. Thus Algorithm˜2 runs in polynomial time. The proof of the lemma is similar to EKKL (20), so here we only give a proof sketch.
@@
According to Theorem˜40 and Lemma 41, we have $T=\Theta(\frac{W({\varepsilon}{U}_{\triangle}+{U}_{\Lambda})}{n\log(n/\delta)\tilde{\ell}_{3}(\hat{G})})=\tilde{O}(W)$. In each inner iteration of Algorithm˜2, note that Algorithm˜3 runs in $O(n^{2})$ time. Denote $\mathbf{X}^{*}$ be the maximizer of the SDP $\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X})$. We use the algorithm of LSW (15) to find an approximate solution of the SDP in Line 6 of Algorithm˜2 in time $\tilde{O}(n^{6}\log^{O(1)}(n/\mu))$, such that, $\left\|(\mathbf{X}^{*})^{-\frac{1}{2}}(\mathbf{X}^{*}-\mathbf{X})(\mathbf{X}^{*})^{-\frac{1}{2}}\right\|_{F}\leq\mu$. Therefore, within $LT$ iterations, Algorithm˜2 can be implemented in time $\tilde{O}(n^{6}W\log^{O(1)}(n))$. According to EKKL (20), the additive error only differs by a constant factor when we choose $\mu=1/n^{O(1)}$. And the privacy loss in each iteration of Algorithm˜2 is still $O(\frac{\ell_{3}(\hat{G})}{\lambda})$, hence the privacy guarantee is the same as before. ∎
@@ 5 The Lower Bound
Given graph $G$, let $\ell_{h}(G)=\max_{(i,j)\in\binom{V}{2}}\sum_{V^{\prime}=(i,j,v_{1},v_{2},\dots,v_{h-2})\in\binom{V}{h}}\prod_{(k,\ell)\in\binom{V^{\prime}}{2}\setminus\{(i,j)\}}\mathbf{w}_{(k,\ell)}$ denotes the local sensitivity of $K_{h}$-motif cuts of $G$. Note that, its local sensitivity is defined as the maximum $K_{h}$-motif cut difference between $G$ and its neighboring graphs. For unweighted graph with maximum degree at most $d$, we have $\ell_{h}(G)=O(d^{h-2})$. For unweighted dense graph, we have $\ell_{h}(G)=\Theta(n^{h-2})$. For unweighted graph generated from $G(n,p)$, we have $\ell_{h}(G)=\Theta(n^{h-2}p^{\frac{h^{2}-h-2}{2}})$ with high probability.
We show the following lower bound. Note that Theorem˜4 follows from the following theorem by setting $h=3$ and $\beta=\frac{3}{4}$.
@@
Note that $\mathbf{A}$ is fixed and does not depend on $G$. Let $\mathbf{x}\in\{0,1\}^{\binom{n}{2}}$ be the indicator vector of $E$. Let $\mathbf{x}_{K_{h}}=f_{K_{h}}(x)\in\{0,1\}^{\binom{n}{h}}$ be the indicator vector of $K_{h}$ in $G$, i.e. for each tuple $i_{1},i_{2},\dots,i_{h}$ of $h$ different indices, $(\mathbf{x}_{K_{h}})_{i_{1},i_{2},\dots,i_{h}}=1$ if the subgraph induced by $i_{1},i_{2},\dots,i_{h}$ is $K_{h}$ and 0 otherwise. Then $\mathbf{A}\mathbf{x}_{K_{h}}$ specifies the $K_{h}$-motif size of all $(S,T)$-cuts in $G$, i.e., we have
[…]
@@
    for each vertex $v\in V$, its degree belongs to the interval $\gamma n/2,2\gamma n]$;
@@
    for each edge $e\in E\cup E^{\prime}$, the number of $K_{h}$-instances containing $e$ belongs to the interval $\frac{\ell_{h}(G)}{2},2\ell_{h}(G)]$;
@@
For any $i\leq h$ and any set $W\subseteq V$, we let $W^{(i)}$ denote the set of all subsets $R\subseteq W$ with exactly $i$ distinct vertices. For any $i\leq h$ and a set $R\subseteq V^{(i)}$, and hih-i other vertices $t_{i+1},\dots,t_{h}$, we let $\chi_{R,t_{i+1},\dots,t_{h}}$ denote the coloring of the potential $K_{h}$ instance induced by the vertex set $R\cup\{t_{i+1},\dots,t_{h}\}$. We define $\chi_{R}=\sum_{t_{i+1},\dots,t_{h}\in V^{(h-i)}}\chi_{R,t_{i+1},\dots,t_{h}}$ and define $\chi_{R,t}=\sum_{t_{i+2},\dots,t_{h}\in V^{(h-i-1)}}\chi_{R,t,t_{i+2},\dots,t_{h}}$. We extend the definition of $\chi$ to any tuple $i_{1},\dots,i_{h}\in V^{(h)}$ by setting $\chi_{i_{1},\dots,i_{h}}=0$ whenever there exists a repeated index in the tuple.
@@
Let $H=(W,E)$ be a graph such that for any fixed $B$ with |B|=i1|B|=i-1, the number of $t\in W$ such that $t$ is connected to all vertices in $B$ is at most $2n\gamma^{i-1}$. Let $S\cup T$ be a random partition of $W$, i.e., each vertex is independently assigned to each class with probability $1/2$. Then
@@
$\geq 2^{-4}\cdot\mathbb{E}\left[\sum_{R\in T_{2}^{(2)}}\left|y_{2}(R)\right|/\sqrt{n\gamma}\right]$
@@
$\Upsilon$ $\leq\mathbb{E}\left[\sum_{s\in T_{1}}\left|\sum_{s_{2}\in S_{2},\dots,s_{h}\in S_{h}}\chi_{s,s_{2},\dots,s_{h}}\right|\right]$
$=\mathbb{E}\left[\sum_{s\in S_{1}^{+}}\left(\sum_{s_{2}\in S_{2},\dots,s_{h}\in S_{h}}\chi_{s,s_{2},\dots,s_{h}}\right)\right]+\mathbb{E}\left[\sum_{s\in S_{1}^{-}}\left(-\sum_{s_{2}\in S_{2},\dots,s_{h}\in S_{h}}\chi_{s,s_{2},\dots,s_{h}}\right)\right]$
$:=\mathbb{E}\left[\left|\chi_{S_{1}^{+},S_{2},\dots,S_{h}}\right|\right]+\cdot\mathbb{E}\left[\left|\chi_{S_{1}^{-},S_{2},\dots,S_{h}}\right|\right]$
@@
$\mathcal{F}=\{\mathcal{J}\subseteq\{2,\dots,h\}:\left|\chi_{\mathcal{J}}\right|\geq(2h)^{-h+\left|\mathcal{J}\right|}\cdot\left|\chi_{S_{1}^{+},S_{2},\dots,S_{h}}\right|\}.$
@@
Let $\mathbf{x}$ be an indicator vector of the edge set of a graph $G=(V,E)$ which satisfies the following properties $(\star)$:
@@
    for each vertex $v\in V$, its degree belongs to the interval $\gamma n/2,2\gamma n]$;
@@
    for each edge $e\in E$, the number of $K_{h}$-instances containing $e$ belongs to the interval $\frac{\ell_{h}(G)}{2},2\ell_{h}(G)]$;
@@
Let $\mathcal{A}$ simply be the algorithm that output an indicator vector $x^{\prime}$ of any graph that satisfies the properties $(\star)$ and that
@@
Part 2) follows from the same argument as those in the proof of Lemma 5.3 in EKKL (20). For part 1), we describe our changes. Formally, we say that $\mathbf{x}\sim X_{\mid Y=\mathbf{y}}$ is good, if $\left\|\mathbf{A}\mathbf{x}_{K_{h}}-\mathbf{y}\right\|_{\infty}\leq\mathrm{disc}_{\mathcal{C}_{\sigma,\gamma}}(\mathbf{A})/2-1$ and the properties $(\star)$ given in the statement of Lemma 50 are satisfied.
Note that the property that $\mathbf{x}\sim X_{\mid Y=\mathbf{y}}$ is good is at least $(1-\frac{1}{\mathrm{poly}(n)})\cdot\beta$. This is true as our assumption that $G\sim G(n,p)$ and that $(\frac{\log n}{n})^{1/(h-1)}\ll p\leq\frac{1}{2}$, it holds that $G$ satisfies the properties $(\star)$ with probability at least $1-\frac{1}{\mathrm{poly}(n)}$. This then finishes the proof of the lemma. ∎
Finally, our lower bound Theorem˜44 follows from the following lemma from EKKL (20), which in turn is built from Lemma 2.1.2 in Bun (16).
@@
Now consider the scaled version of $G$ described in Theorem˜44, where the weight vector $\mathbf{w}$ of $G$ is scaled by a factor of $\frac{1}{\varepsilon}$, resulting in a new weight vector $\frac{1}{\varepsilon}\cdot\mathbf{w}$. Note that the total weight is $W=\frac{m}{\varepsilon}$. When we let $\overline{T}_{\mathbf{w}}=\sqrt{Wn}\cdot\ell_{h}(G)$, then $\overline{T}_{\frac{1}{\varepsilon}\cdot\mathbf{w}}=\frac{1}{\varepsilon^{\binom{h}{2}-\frac{1}{2}}}\overline{T}_{\mathbf{w}}$. By a similar analysis to Lemma 53, we can establish the bound
@@ 6 Conclusion
Acknowledgements
We thank Jalaj Upadhyay for helpful discussions and the anonymous reviewers for their valuable comments on an earlier version of this paper.
@@ A.1 Synthetic Graphs with Multiplicative Errors and Interactive Solutions
Finally, we note that an interactive solution for preserving privacy in the motif cut structure could also be considered (see Section˜1.2 for considerations related to edge cuts). In this scenario, data analysts could specify any cut $(S,V\setminus S)$ with the aim of determining (with some acceptable error) the number of triangles or motifs connecting the two groups while maintaining privacy. However, our primary focus is on a stronger, non-interactive solution, i.e. to release a private synthetic dataset: a new, private graph that approximately preserves the motif cut function of the original graph.
@@ A.2 Some Tempting Approaches That Do Not Work
The third approach is to make use of the triangle-motif-weighted graph associated with the input graph [6]. That is, one can first convert the original graph G{G} into a triangle-motif weighted graph ${G}_{\triangle}$ with a weight vector $\mathbf{{w}}_{\triangle}$, where $\mathbf{{w}}_{\triangle}(e)$ denotes the sum of the weights of triangles containing the endpoints of edge $e$ simultaneously. For triangle-motif weighted graphs, there is a useful property: the size of the edge cut $(S,V\setminus S)$ in ${G}_{\triangle}$ is exactly twice the sum of the weights of triangles crossing the cut in $G$. Therefore, a naive approach would be to use a private edge cut release algorithm, such as the one proposed in [37], on the triangle-motif weighted graph. However, after applying the DP algorithm on ${G}_{\triangle}$, the resulting graph $H$ may not correspond to a triangle-motif weighted graph, i.e. there does not exist a graph whose triangle-motif weighted graph is $H$. As a result, through this approach, we can only privately release the values of triangle-motif cuts but cannot release a synthetic graph.
Appendix B Deferred Preprocessing Step in Section˜3
@@
    Set $u_{\max}=\max_{e\in\binom{V}{2}}\mathbf{u}_{e}$. Recall that $\ell_{3}(\hat{G})=\max_{(i,j)\in\binom{V}{2}}\sum_{s\in V\setminus\{i,j\}}\mathbf{\hat{w}}_{(i,s)}\mathbf{\hat{w}}_{(j,s)}$ denotes the local sensitivity of triangle-motif cuts of $\hat{G}$. We further define $\tilde{\ell}_{3}(\hat{G})=\ell_{3}(\hat{G})+u_{\max}(\mathrm{Lap}(1/\varepsilon_{3})+\log(6n^{2}/\beta)/\varepsilon_{3})$.
[…]
@@
Assume that $\hat{W}=\Omega(\frac{1/\beta}{\varepsilon})$ and $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$. We refer to this as the non-degenerate case. (When this assumption does not hold, we call it the degenerate case. We will provide a case analysis in Corollary 42.) Recall that $\ell_{3}(\overline{G})=\max_{(i,j)\in\binom{V}{2}}\sum_{s\in V\setminus\{i,j\}}\overline{\mathbf{w}}_{(i,s)}\overline{\mathbf{w}}_{(j,s)}$ is the local sensitivity of triangle cuts in $\overline{G}$. Since $\hat{W}=\Omega(\frac{1/\beta}{\varepsilon})$, we have $)\ell_{3}(\overline{G})=(\frac{\overline{W}}{\hat{W}})^{2}\ell_{3}(\hat{G})=\Theta(\ell_{3}(\hat{G}$)). Additionally, by the assumption $\ell_{3}(\hat{G})=\Omega(\frac{w_{\max}\log^{2}(n/\beta)}{\varepsilon^{2}})$ and the fact $\tilde{\ell}_{3}(\hat{G})\leq\ell_{3}(\hat{G})+2u_{\max}\log(6n^{2}/\beta)/\varepsilon_{3}\leq\ell_{3}(\hat{G})+O(w_{\max}\log^{2}(n/\beta)/\varepsilon_{2}\varepsilon_{3})$, $\varepsilon_{2}=\varepsilon_{3}=\varepsilon/6$, it holds that $)\tilde{\ell}_{3}(\hat{G})=\Theta(\ell_{3}(\hat{G}$)).